A competitive, cost-minimizing firm has the production function f (x, y) = min(x,2y) and uses positive amounts of both inputs. If the price of x doubles and the price of y triples, then the cost of production will exactly double.

The answer is False. Plz show me exactly how to solve this.
Answer #1

As this is a fixed proportions production function, so two inputs are always used in given proportion:
x = 2y
Let price of x be Px and price of y be Py.
Cost, C = x*Px + y*Py = 2y*Px + y*Py = (2Px + Py)y

New price of x is 2Px and price of y is 3Py.
New cost, C' = x*(2Px) + y*(3Py) = 2y*(2Px) + y*(3Py) = (4Px + 3Py)y

Double of C = 2C = 2*(2Px + Py)y = (4Px + 2Py)y
We can see that C' does not equal C. So, cost of production will not exactly double.